# Hypergraph-MLP: Learning on Hypergraphs without Message Passing

### 15 Dec 2023 v1

@@ 1 Introduction
Fig. 1: The training paradigm of the Hypergraph-MLP and existing hypergraph neural networks for hypergraph node classification, where $\mathbf{X}_{\mathcal{V}}$ denotes node features, $\mathbf{H}$ is the incidence matrix, $\mathbf{Y}_{\mathcal{V}}$ denotes node labels, and only partial labels in $\mathbf{Y}_{\mathcal{V}}$ are used at training. Key components of the Hypergraph-MLP are highlighted in red.
@@ 2 Related Works
Learning on hypergraphs. Processing hypergraph-structured data efficiently is important in both machine learning and signal processing. Many studies focus on designing neural networks to process these data. Recent research [13] shows that most hypergraph neural networks [14, 15, 16, 17] follow a two-step message passing paradigm: node features are first aggregated to hyperedges to update hyperedge embeddings, which are aggregated back to nodes to update node embeddings. These message-passing-based models have several limitations, including oversmoothing [7, 8], high inference latency [6, 9], and sensitivity to structural perturbations [10, 11]. To tackle this challenge, we propose Hypergraph-MLP for processing hypergraph-structured data without message passing.
Multilayer perceptions for structural data. In the literature, many recent works aim to create more efficient and robust models for processing data on graphs using MLPs. These models incorporate the graph structure as part of the supervision during training, either through the GNN-MLP knowledge distillation loss [9, 18] or the neighbour contrastive loss [19]. These studies improve MLPs by explicitly considering pairwise node relations, which makes them not well-suited for hypergraph-structured data with higher-order node relations. To our knowledge, our work is the first to demonstrate the effectiveness of MLPs in processing hypergraph-structured data.
@@ 4.1 MLP-based Model
Inspired by previous works in graph machine learning [9, 19, 18], we design each layer of the MLP-based model as:
@@
where $D(\cdot)$ is the dropout function, $LN(\cdot)$ is the layer normalization function, $\sigma(\cdot)$ is the activation function, and $\mathbf{\Theta}^{(l)}\in\mathbb{R}^{d^{(l-1)}\times d^{(l)}}$ denotes the learnable parameters for layer $l$. For the node classification task, we add a classifier to the output of the MLP-based model to generate the node label logits as:
@@ 4.2 Hypergraph-Smoothness-based Loss Function
Hypergraph node embedding smoothness prior. Taking the inspiration from a recent work [12], we propose a smoothness prior on the hypergraph node embedding: embeddings of nodes in a hyperedge are highly correlated via the hyperedge connecting these nodes. Under this prior, we capture the relation between nodes and hyperedges in the latent space by an incidence graph [20, 21]. For the hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, we construct the unique incidence graph as a bipartite graph $\mathcal{G}=\{\mathcal{V}\bigcup\mathcal{V}^{\prime},\mathcal{E}_{\mathcal{G}},\mathbf{L}_{\mathbf{H}}\}$. Here $v_{i}\in\mathcal{V}$ is node $v_{i}$ in $\mathcal{H}$, $v_{e_{j}}\in\mathcal{V}^{\prime}$ corresponds to the hyperedge $e_{j}$ in $\mathcal{H}$. An edge exists between $v_{i}$ and $v_{e_{j}}$ if and only if $e_{j}$ contains $v_{i}$ in $\mathcal{H}$. The structure of $\mathcal{G}$ is represented as a graph Laplacian matrix:
@@
Loss function via maximum likelihood estimation. Our goal is to create a loss function derived from the likelihood of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$, thus allowing the MLP-based model to utilize the hypergraph structural information by generating node embeddings that follow the smoothness prior mentioned above. With the distribution in Eq. (4) we have $p(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}}|\mathbf{H})\!\propto\!e^{-\frac{1}{2}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]\mathbf{L}_{\mathbf{H}}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}}$. Omitting the constant term, the negative log likelihood function of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$ is:
$\mathcal{L}(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}}|\mathbf{H})$ $\propto$ $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]\mathbf{L}_{\mathbf{H}}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$ (5)
@@ 4.3 Training & Inference
For node classification, we train the MLP-based model using a combination of the previously mentioned hypergraph-smoothness-based loss function and the cross-entropy loss function for the node classification task:
@@ 4.4 Discussion
Cora Citeseer Pubmed DBLP 20Newsgroups NTU2012 House Avg Mean
@@
Cora Citeseer Pubmed DBLP 20Newsgroups NTU2012 House Avg Mean
@@ 5.1 Experimental Settings
Datasets. We use seven publicly available datasets, including academic hypergraphs (Cora, Citeseer, Pubmed, and DBLP), adapted from [14], 20Newsgroups from UCI’s Categorical Machine Learning Repository [26], NTU2012 from computer vision [27], and House from politics [28]. For the House dataset, lacking node features, we follow [13] and use Gaussian random vectors instead, where the standard deviation of the added Gaussian features is set as $0.6$. Details of the datasets used are in Table 1.
@@
Our implementation. Following a previous work [13], we randomly partitioned the dataset into training, validation, and test sets using a 50%/25%/25% split. All experiments were performed on an RTX 3090 using PyTorch. Our code is available at: https://github.com/tbh-98/Hypergraph-MLP.
@@ 5.2 Results & Analysis
Performance comparison on clean datasets. The results for real-world hypergraph node classification benchmarks, conducted without considering structural perturbations, are summarized in Table 2 and Table 3. In Table 2, we present the ACC comparison of the proposed Hypergraph-MLP against all baseline methods. Our results demonstrate that Hypergraph-MLP not only outperforms the baseline methods across four datasets (Cora, Citeseer, 20Newsgroups, and House) but also achieves the highest average mean ACC over all of the datasets. Furthermore, Hypergraph-MLP consistently outperforms the standard MLP across all the datasets. These findings underscore the efficacy of the hypergraph-smoothness-based loss function in effectively enabling the MLP-based model to leverage hypergraph structural information, all while not explicitly using the structure itself as input for message passing. As illustrated in Table 3, Hypergraph-MLP demonstrates the fastest inference speed on all the seven chosen datasets compared to existing message-passing-based hypergraph neural networks. Notably, across the seven datasets, Hypergraph-MLP achieves an average mean inference time that is only $49\%$ of the time required by the fastest message-passing-based model (HyperGCN) and a mere $5\%$ of the time needed by the most time-consuming message-passing-based model (AllDeepSets). This attribute renders Hypergraph-MLP well-suited for applications with stringent latency constraints, such as autonomous driving and recommender systems, in contrast to previous hypergraph neural networks.

### 19 Mar 2024 v2

@@ 1 Introduction
Fig. 1: The training paradigm of the Hypergraph-MLP and hypergraph neural networks for hypergraph node classification, where $\mathbf{X}_{\mathcal{V}}$ denotes node features, $\mathbf{H}$ is the incidence matrix, $\mathbf{Y}_{\mathcal{V}}$ denotes node labels, and only partial labels in $\mathbf{Y}_{\mathcal{V}}$ are used at training.
@@ 2 Related Works
Learning on hypergraphs. There are many useful tools for modelling data with higher-order interactions, e.g., hypergraphs, and simplicial complex [13, 14, 15]. This paper focuses on the process of hypergraph-structured data. Recent research [16] shows that most neural networks designed for hypergraphs follow a two-stage message passing paradigm: node features are first aggregated to hyperedges to update hyperedge embeddings, which are aggregated back to nodes to update node embeddings. These message-passing-based models have several limitations, including oversmoothing [7], high inference latency [6, 9], and sensitivity to structural perturbations [10, 11]. For this challenge, we propose Hypergraph-MLP for processing hypergraph-structured data without message passing.
Multilayer perceptions for structural data. In the literature, many recent works aim to create more efficient and robust models for processing data on graphs using MLPs. These models incorporate the graph structure as part of the supervision during training, either through the GNN-MLP knowledge distillation loss [9] or the neighbour contrastive loss [17]. These studies improve MLPs by explicitly considering pairwise node relations, which makes them not well-suited for hypergraph-structured data with higher-order node relations. To our knowledge, our work is the first to demonstrate the effectiveness of MLPs in processing hypergraph-structured data.
@@ 4.1 MLP-based Model
Inspired by previous works in graph machine learning [9, 17], we design each layer of the MLP-based model as:
@@
where $D(\cdot)$ is the dropout function, $LN(\cdot)$ is the layer normalization function, $\sigma(\cdot)$ is the activation function, and $\mathbf{\Theta}^{(l)}\in\mathbb{R}^{d^{(l-1)}\times d^{(l)}}$ denotes the learnable parameters for layer $l$. For the node classification task, we generate the node label logits by:
@@ 4.2 Hypergraph-Smoothness-based Loss Function
Hypergraph node embedding smoothness prior. Taking the inspiration from a recent work [12], we propose a smoothness prior on the hypergraph node embedding: embeddings of nodes in a hyperedge are highly correlated via the hyperedge connecting these nodes. Under this prior, we capture the relation between nodes and hyperedges in the latent space by an incidence graph [21]. For the hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, we construct the unique incidence graph as a bipartite graph $\mathcal{G}=\{\mathcal{V}\bigcup\mathcal{V}^{\prime},\mathcal{E}_{\mathcal{G}},\mathbf{L}_{\mathbf{H}}\}$. Here $v_{i}\in\mathcal{V}$ is node $v_{i}$ in $\mathcal{H}$, $v_{e_{j}}\in\mathcal{V}^{\prime}$ corresponds to the hyperedge $e_{j}$ in $\mathcal{H}$. An edge exists between $v_{i}$ and $v_{e_{j}}$ if and only if $e_{j}$ contains $v_{i}$ in $\mathcal{H}$. The structure of $\mathcal{G}$ is represented as a graph Laplacian matrix:
@@
Loss function via maximum likelihood estimation. Our goal is to create a loss function derived from the likelihood of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$, thus allowing the MLP-based model to utilize the hypergraph structural information by generating node embeddings that follow the smoothness prior mentioned above. With the distribution in Eq. (4) we have $p(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}},\mathbf{H})\!\propto\!e^{-\frac{1}{2}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]\mathbf{L}_{\mathbf{H}}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}}$. Omitting the constant term, the negative log likelihood function of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$ is:
${\cal L}(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}}|\mathbf{H})$ $\propto$ $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]\mathbf{L}_{\mathbf{H}}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$ (5)
@@ 4.3 Training & Inference
For node classification, we train the MLP-based model using a combination of the aforementioned hypergraph-smoothness-based loss function and the cross-entropy loss function for node classification:
@@ 4.4 Discussion
Cora Citeseer Pubmed DBLP 20News NTU2012 House Avg Mean
@@
Cora Citeseer Pubmed DBLP 20News NTU2012 House Avg Mean
@@ 5.1 Experimental Settings
Datasets. We use seven public datasets, including academic hypergraphs (Cora, Citeseer, Pubmed, and DBLP), adapted from [20], 20News from UCI’s Categorical Machine Learning Repository [26], NTU2012 from computer vision [27], and House from politics [28]. For the House dataset, lacking node features, we follow [16] and use Gaussian random vectors instead, where the standard deviation of the added Gaussian features is set as $0.6$. Notably, the datasets 20News and House represent examples of heterophilic hypergraphs, while Cora, Citeseer, Pubmed, DBLP, and NTU2012 exemplify homophilic hypergraphs. Details of the datasets used are in Table 1.
@@
Our implementation. Following a previous work [16], we randomly partitioned the dataset into training, validation, and test sets using a 50%/25%/25% split. All experiments were performed on an RTX 3090 using PyTorch. We identified the optimal value of $\alpha$ in Eq. (8) through grid search. Our code is available at: https://github.com/tbh-98/Hypergraph-MLP.
@@ 5.2 Results & Analysis
Performance comparison on clean datasets. The results for real-world hypergraph node classification benchmarks, conducted without considering structural perturbations, are summarized in Table 2 and Table 3. In Table 2, we present the ACC comparison of the proposed Hypergraph-MLP against all baseline methods. Our results demonstrate that Hypergraph-MLP not only outperforms the baseline methods across four datasets (Cora, Citeseer, 20News, and House) but also achieves the highest average mean ACC over all of the datasets. Furthermore, Hypergraph-MLP consistently outperforms the standard MLP across all the datasets. These findings underscore the efficacy of the hypergraph-smoothness-based loss function in effectively enabling the MLP-based model to leverage hypergraph structural information, all while not explicitly using the structure itself as input for message passing. As illustrated in Table 3, Hypergraph-MLP demonstrates the fastest inference speed on all the seven chosen datasets compared to existing message-passing-based hypergraph neural networks. Notably, across the seven datasets, Hypergraph-MLP achieves an average mean inference time that is only $49\%$ of the time required by the fastest message-passing-based model (HyperGCN) and a mere $5\%$ of the time needed by the most time-consuming message-passing-based model (AllDeepSets). This attribute renders Hypergraph-MLP well-suited for applications with stringent latency constraints, such as autonomous driving and recommender systems, in contrast to previous hypergraph neural networks.
@@ 6 Conclusion
Limitation and future work. The Hypergraph-MLP is currently limited to transductive learning, with the hyperparameter $\alpha$ in Eq. (8) can only be optimized through grid search. The potential future works include extending to the inductive learning setting and improving the hyperparameter selection mechanism.
Acknowledgment
[…]